# Unveiling Impact of Frequency Components on Membership Inference Attacks for Diffusion Models

### v1

@@ 5.1 Experiment Setup
Datasets and Models. We adhere to the stringent assumption that both member and hold-out data reside within the same distribution. For DDIM, we used the STL10-U [7], CIFAR-100 [17], and Tiny-ImageNet(Tiny-IN) [7] datasets for training, respectively. Specifically, we randomly selected 50% of the training set to serve as member data, while the remaining 50% was designated as hold-out data. For text-to-image diffusion models, we employed 416/417 samples on Pokémon [18], 2500/2500 samples on MS-COCO [23], and 1000/1000 samples on Flickr [43] as the member/hold-out dataset, utilizing stable diffusion v1-4 [6] for fine-tuning. Furthermore, for pre-trained diffusion models, we selected stable diffusion v1-4 and v1-5 [29] as attack targets. We adhere to the settings in [9, 44], employing Laion-MI [9] dataset to ensure that both member and hold-out data exhibit the same distribution. Detailed training information is provided in Appendix E.2.
Evaluation Metrics. We use the established metrics employed in prior research [24, 8, 16], which include the Attack Success Rate (ASR), the Area Under the Curve (AUC), and the True Positive Rate (TPR) at 1% False Positive Rate (FPR) (denoted as TPR@1%FPR).
Baselines. We employed the SecMI [8], PIA [16], and Naive [24] as our baselines for comparison. These approaches are characterized as query-based attacks and operate effectively within a grey-box setting. We adopt the parameters advised in their respective publications.
@@
Method STL10-U CIFAR-100 Tiny-IN
@@
Method Pokémon MS-COCO Flickr
@@ 5.2 Overall Performance
Stable Diffusion Models. The experimental results for the fine-tuned stable diffusion attacks are presented in Tab. 4. Based on the analysis of the average improvement over baselines, we observed that the ASR improved by 10.01%, AUC improved by 9.01%, and the improvement in the TPR@1% FPR reached as high as 19.73% on the MS-COCO dataset. Notably, on Flickr, our method achieved the highest TPR@1% FPR improvement of 51.01% in Naive. These results strongly indicate that our method significantly enhances the attack efficacy of the baselines across diverse data distributions and scales by mitigating high-frequency deficiency. Moreover, we have conducted tests on the pre-trained stable diffusion, and the results indicate that the filter exhibits only a modest effect, possibly due to the inherent shortcomings of the baselines. A detailed discussion is provided in Appendix E.3.
@@ 5.3 In-depth Analysis of Attack Performance
We have theoretically proven that removing high-frequency deficiency can amplify the distinction between member and hold-out data. To further validate our conjectures, we visualize the membership scores of the baselines before and after applying the filter. As illustrated in Fig. 2, it is evident that the distribution gap between member data and hold-out data has increased noticeably after applying our method. Blue boxes mark the areas where member and hold-out data interleave, indicating member/hold-out data are indistinguishable by thresholds. After applying the filter, we observe a significant reduction in sample interleavements. This compelling evidence validates our conjectures and demonstrates the filter’s effectiveness. More visualizations will be presented in Appendix E.4.
@@
5.4 Robustness Analysis
Weaker Overfitting. When fine-tuning stable diffusion, the number of iterations is typically determined by the user’s specific needs. In some scenarios, extensive iterations are not required to fine-tune the model. To simulate a model with a lower degree of overfitting, we halved the number of iterations across all datasets, as referenced in [44]. As shown in Tab. 5, the filter demonstrates nice effectiveness. Taking the Flickr dataset as an example, after applying the filter, the ASR improved by an average of 4.34%, with a maximum improvement of 9.25% in Naive. The AUC increased by an average of 7.04%, with the Naive demonstrating the most significant improvement of 14.53%. Compared to the default settings, the filter’s effect is diminished, which is directly linked to the performance decline of the baseline under the weaker overfitting. Therefore, the results can still prove the effectiveness of the high-frequency filter.
@@ 5.4 Robustness Analysis
Method Pokémon MS-COCO Flickr
@@
5.5 Ablation Study
To investigate the impact of the filter under different hyperparameter settings, we adjusted the high-frequency threshold $r_{t}$ and the filtering parameter $s$. Experiments were conducted with various values for $r_{t}$ and $s$ on the MS-COCO dataset using Naive attack, and the results are presented in Tab. 6. From the experimental analysis, we recommend a value range for $r_{t}$ of $[3,10]$ and for $s$ of $[0.0,0.3]$. Within this range, the filter achieves optimal performance, significantly enhancing the baseline performance while exhibiting low sensitivity to changes in hyperparameters, further demonstrating the robustness of the filter. When $r_{t}=1$, the high-frequency threshold is set too low, leading to most frequency components of the image being filtered. This contradicts our intention to only suppress high-frequency deficiency, resulting in a significant decline in attack performance. When $s=0.4$ and $s=0.5$, the filtering effect on the high-frequency components is weakened. Although the baseline performance improves significantly, it does not reach the optimal state. More ablation experiments will be presented in Appendix E.5.
$s/r_{t}$ $s=0.0$ $s=0.1$ $s=0.2$ $s=0.3$ $s=0.4$ $s=0.5$
ASR AUC ASR AUC ASR AUC ASR AUC ASR AUC ASR AUC
5.6 Impact of Data Augmentation
In the training process, data augmentation techniques are typically employed by default to prevent overfitting. For example, DDIM uses RandomHorizontalFlip during training, and stable diffusion employs Random-Crop and Random-Flip by default during fine-tuning [10]. Based on this context, this study refers to the experimental strategies regarding the impact of data augmentation on MIAs from [44, 8], eliminating data augmentation from the training process and attacking the trained models. As indicated in Tab. 7, removing data augmentation while fine-tuning the stable diffusion enhances the attack performance of the baselines, and the filter still shows excellent improvement. Notably, this improvement is particularly pronounced for Naive, where the TPR@1% FPR increased by 40.01%, 54.81%, and 58.70% across different datasets, while the AUC reached 96.81%, 98.42%, and 98.21%, respectively. Moreover, other baselines also exhibited significant performance improvements. Detailed analysis of the data augmentation in DDIM will be provided in Appendix E.7.
Table 7: After removing data augmentation, the attack performance in fine-tuned stable diffusion.
Method Pokémon MS-COCO Flickr
ASR AUC TPR@1%FPR ASR AUC TPR@1%FPR ASR AUC TPR@1%FPR
Naive 80.00 86.45 5.99 80.09 87.61 5.99 79.50 86.17 12.80
Naive+F 92.00 96.81 46.00 95.60 98.42 60.80 93.50 98.21 71.50
SecMI 81.08 86.67 5.30 80.90 88.23 11.20 80.69 87.55 14.80
SecMI+F 85.00 90.42 23.00 86.50 92.61 13.60 85.50 91.76 35.50
PIA 75.49 82.64 10.60 69.90 74.36 6.40 68.40 71.68 7.00
PIA+F 87.00 92.54 28.00 79.10 86.37 25.00 76.50 82.44 31.00
Avg+ +9.14 +8.00 +25.03 +10.10 +9.07 +25.27 +8.97 +9.00 +34.47
@@ 6 Conclusion
In this paper, we define a general paradigm for the pixel-wise error-based MIAs for diffusion models. Under this general paradigm, we find that the current attacks ignore the intrinsic deficiency of the diffusion model in handling the high-frequency components, which results in limited attack performance. To address this, we introduce a simple and efficient method which mitigates the negative impact of high-frequency deficiency on MIAs by filtering images’ high-frequency information. Experimental results reveal that our method can be seamlessly incorporated into attacks within the general paradigm, significantly enhancing attack performance across diverse settings.
@@
Appendix A Denoising Diffusion Probabilistic Model (DDPM)
DDPM [13] consists of a forward process and a denoising process. In the forward process, DDPM transitions from an intractable data distribution, represented as $x_{0}\sim q_{0}(x_{0})$, to a Gaussian distribution $q_{T}(x_{T})\sim\mathcal{N}(x_{T};0,I)$. This transition is achieved by progressively adding Gaussian noise to the original image $x_{0}$. Consequently, the transition distribution at timestep $t$ is defined as follows:
@@ Appendix A Denoising Diffusion Probabilistic Model (DDPM)
where $\epsilon_{\theta}(x_{t},t)$ is predicted by the diffusion models. The denoising (or reverse) process shares the same functional form as the forward process [34]. It is expressed as a Gaussian transition characterized by a learned mean and a fixed variance:
@@ Appendix B General Paradigm
Naive [24] determines member data through the loss function, specifically judging based on the distance between the added noise and the predicted noise. SecMI [8] leverages diffusion inversion to obtain the ground truth at step $t$ and step $t-1$. Based on this, the model predicts $x_{t-1}$ from the ground truth at step $t$ and assesses whether a sample is a member by examining the distance between the predicted $x_{t-1}$ and its ground truth. PIA [16] retrieves an initial noise through a proximal initialization process, then uses the model to predict noise for samples containing that initial noise. Finally, it evaluates the membership status based on the distance between the predicted noise and the initial noise. In this section, we will demonstrate that the baselines Naive, PIA and SecMI can be translated into the general paradigm we have defined, and here we use the $\ell_{1}$ norm as an example.
@@
$||\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\,t\right)||<c,$ (B.2)
@@
$||\sqrt{1-\bar{\alpha}_{t}}\epsilon-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\,t\right)||<\sqrt{1-\bar{\alpha}_{t}}c,$ (B.3)
@@
$||\sqrt{1-\bar{\alpha}_{t}}\epsilon-x_{t}+x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\,t\right)||<\sqrt{1-\bar{\alpha}_{t}}c.$ (B.4)
@@
$||\frac{\sqrt{1-\bar{\alpha}_{t}}\epsilon-x_{t}+x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\,t\right)}{\sqrt{\bar{\alpha_{t}}}}||<\frac{\sqrt{1-\bar{\alpha}_{t}}c}{\sqrt{\bar{\alpha_{t}}}}.$ (B.5)
@@
$||\frac{\sqrt{1-\bar{\alpha}_{t}}\epsilon(x_{0},0)-x_{t}+x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon(x_{0},0)\,,\,t\right)}{\sqrt{\bar{\alpha_{t}}}}||<\frac{\sqrt{1-\bar{\alpha}_{t}}c}{\sqrt{\bar{\alpha_{t}}}}.$ (B.9)
@@
SecMI: Inspired by recent works on deterministic reversing and sampling from diffusion models [37, 15], SecMI used DDIM and DDIM inversion deterministic sampling in the forward and backward processes for the samples to be tested:
@@
Denote by $\boldsymbol{\Phi}_{\theta}(x_{s},t)$ the deterministic reverse, i.e., from $x_{s}$ to $x_{t}$ ($s<t$):
$x_{t}=\boldsymbol{\Phi}_{\theta}(x_{s},t)=\phi_{\theta}(\cdots\phi_{\theta}(\phi_{\theta}(x_{s},s),s+1),t-1),$ (B.14)
and $\boldsymbol{\Psi}_{\theta}(x_{t},s)$ the deterministic denoise process, i.e., from $x_{t}$ to $x_{s}$:
$x_{s}=\boldsymbol{\Psi}_{\theta}(x_{t},s)=\psi_{\theta}(\cdots\psi_{\theta}(\psi_{\theta}(x_{t},t),t-1),s+1).$ (B.15)
@@
where $\tilde{x}_{t}=\boldsymbol{\Phi}_{\theta}(x_{0},t)$. A series of transformations are designed to fetch $x_{t+1}^{target}$ and $x_{t}^{target}$ and predict $x_{t}$ by $x_{t+1}^{target}$. Therefore, Eq. B.16 can be expressed as:
$||x_{t}-x_{t}^{target}||\leq\tau,$ (B.17)
@@
Appendix C Proof of Theorem 1
Theorem 1.
Assume membership scores $s$ follow a normal distribution. Let membership scores of member data $s_{1}\sim N(0,\sigma_{M}^{2})$ and hold-out data $s_{2}\sim N(0,\sigma_{H}^{2})$, where $\sigma_{H}>\sigma_{M}$. The membership advantage of algorithm $\mathcal{A}$ is:
$Adv^{M}(\mathcal{A})=erf\left(\frac{\sigma_{H}}{\sigma_{M}}\sqrt{\frac{\ln(\sigma_{H}/\sigma_{M})}{(\sigma_{H}/\sigma_{M})^{2}-1}}\right)-erf\left(\sqrt{\frac{\ln(\sigma_{H}/\sigma_{M})}{(\sigma_{H}/\sigma_{M})^{2}-1}}\right).$ (C.1)

### v2

@@ 5.1 Experiment Setup
Datasets and Models. We adhere to the stringent assumption that both member and hold-out data reside within the same distribution. For DDIM, we used the STL10-U (Coates et al., 2011), CIFAR-100 (Krizhevsky et al., 2009), and Tiny-ImageNet(Tiny-IN) (Deng et al., 2009) datasets for training, respectively. Specifically, we randomly selected 50% of the training set to serve as member data, while the remaining 50% was designated as hold-out data. For text-to-image diffusion models, we employed 416/417 samples on Pokémon (Lambda, 2023), 2500/2500 samples on MS-COCO (Lin et al., 2014), and 1000/1000 samples on Flickr (Young et al., 2014) as the member/hold-out dataset, utilizing stable diffusion v1-4 (CompVis, 2024) for fine-tuning. Furthermore, for pre-trained diffusion models, we selected stable diffusion v1-4 and v1-5 (RunwayML, 2024) as attack targets. We adhere to the settings in (Dubiński et al., 2024; Zhai et al., 2024), employing Laion-MI (Dubiński et al., 2024) dataset to ensure that both member and hold-out data exhibit the same distribution. Detailed training information is provided in Appendix D.3.
Evaluation Metrics. We use the established metrics employed in prior research (Matsumoto et al., 2023; Duan et al., 2023; Kong et al., 2023), which include the Attack Success Rate (ASR), the Area Under the Curve (AUC), and the True Positive Rate (TPR) at 1% False Positive Rate (FPR) (denoted as TPR@1%FPR).
Baselines. We employed the SecMI (Duan et al., 2023), PIA (Kong et al., 2023), and Naive (Matsumoto et al., 2023) as our baselines for comparison. These approaches are characterized as error-based attacks and operate effectively within a grey-box setting. We adopt the parameters advised in their respective publications.
@@
Method
STL10-U CIFAR-100 Tiny-IN
@@
Method
Pokémon MS-COCO Flickr
@@ 5.2 Overall Performance
Stable Diffusion Models. The experimental results for the fine-tuned stable diffusion attacks are presented in Tab. 3. Based on the analysis of the average improvement over baselines, we observed that the ASR improved by 10.01%, AUC improved by 9.01%, and the improvement in the TPR@1% FPR reached as high as 19.73% on the MS-COCO dataset. Notably, on Flickr, our method achieved the highest TPR@1% FPR improvement of 51.01% in Naive. These results strongly indicate that our method significantly enhances the attack efficacy of the baselines across diverse data distributions and scales by mitigating high-frequency deficiency. We emphasize the generality of our proposed method and validate its effectiveness in other classes of attack methods in Appendix D.5. Moreover, we have conducted tests on the pre-trained stable diffusion, and the results indicate that the filter exhibits only a modest effect, possibly due to the inherent shortcomings of the baselines. A detailed discussion is provided in Appendix D.4.
@@ 5.3 In-depth Analysis of Attack Performance
We have theoretically proven that removing high-frequency deficiency can amplify the distinction between member and hold-out data. To further validate our conjectures, we visualize the membership scores of the baselines before and after applying the filter. As illustrated in Fig. 2, it is evident that the distribution gap between member data and hold-out data has increased noticeably after applying our method. Blue boxes mark the areas where member and hold-out data interleave, indicating member/hold-out data are indistinguishable by thresholds. After applying the filter, we observe a significant reduction in sample interleavements. This compelling evidence validates our conjectures and demonstrates the filter’s effectiveness. More visualizations will be presented in Appendix D.6.
@@
5.4 Ablation Study
To investigate the impact of the filter under different hyperparameter settings, we adjusted the high-frequency threshold $r_{t}$ and the filtering parameter $s$. Experiments were conducted with various values for $r_{t}$ and $s$ on the MS-COCO dataset using Naive attack, and the results are presented in Tab. 4. From the experimental analysis, we recommend a value range for $r_{t}$ of $[3,10]$ and for $s$ of $[0.0,0.3]$. Within this range, the filter achieves optimal performance, significantly enhancing the baseline performance while exhibiting low sensitivity to changes in hyperparameters, further demonstrating the robustness of the filter. When $r_{t}=1$, the high-frequency threshold is set too low, leading to most frequency components of the image being filtered. This contradicts our intention to only suppress high-frequency deficiency, resulting in a significant decline in attack performance. When $s=0.4$ and $s=0.5$, the filtering effect on the high-frequency components is weakened. Although the baseline performance improves significantly, it does not reach the optimal state. More ablation experiments will be presented in Appendix D.8.
$s/r_{t}$
$s=0.0$ $s=0.1$ $s=0.2$ $s=0.3$ $s=0.4$ $s=0.5$
ASR AUC ASR AUC ASR AUC ASR AUC ASR AUC ASR AUC
5.5 more stringent attack conditions
Weaker Overfitting. When fine-tuning stable diffusion, the number of iterations is typically determined by the user’s specific needs. In some scenarios, extensive iterations are not required to fine-tune the model. To simulate a model with a lower degree of overfitting, we halved the number of iterations across all datasets, as referenced in Zhai et al. (2024). As shown in Tab. 5, the filter demonstrates nice effectiveness. Taking the Flickr dataset as an example, after applying the filter, the ASR improved by an average of 4.34%, with a maximum improvement of 9.25% in Naive. The AUC increased by an average of 7.04%, with the Naive demonstrating the most significant improvement of 14.53%. Compared to the default settings, the filter’s effect is diminished, which is directly linked to the performance decline of the baseline under the weaker overfitting. Therefore, the results can still prove the effectiveness of the high-frequency filter.
@@ 5.4 Ablation Study
Method
Pokémon MS-COCO Flickr
@@
Without Captions. We investigated the effectiveness of our method when the attacker has no access to the corresponding text of the images. The results demonstrate that our method remains effective under this condition. Detailed experiments and results are provided in the Appendix D.7.
5.6 Impact of Defense
To investigate the impact of defense mechanisms on our method, we examine two defenses against membership inference attacks (MIAs): the commonly used data augmentation strategy and the state-of-the-art memory mitigation technique $SS_{e_{i}}$ (Wen et al., 2024). Recall that MIAs primarily benefit from overfitting. Data augmentation techniques are typically employed to prevent overfitting. During the fine-tuning process of stable diffusion, it employs Random-Crop and Random-Flip by default (Face, 2024). Memory mitigation dynamically evaluates the model’s memory of the data during training and adjusts the training process accordingly, thereby significantly reducing overfitting and training data leakage. Both serve as effective defenses against MIAs. As shown in Tab. 6, the performance of the baseline methods gradually decreases with the inclusion of defenses. Nevertheless, our method continues to exhibit strong performance, further confirming its robustness in the presence of such defenses. More detailed results are provided in the Appendix D.11.
Table 6: Attack performance AUC under the defenses. With the introduction of defense mechanisms, the baselines exhibit varying degrees of performance degradation. Our method remains effective and continues to deliver a clear performance improvement over the baselines under these defenses.
$SS_{e_{i}}$ DataAug Naive Naive+F Gain SecMI SecMI+F Gain PIA PIA+F Gain
$\times$ $\times$ 87.95 96.81 +8.86 86.67 91.42 +4.75 82.64 92.54 +9.90
$\times$ $\checkmark$ 86.87 94.14 +7.27 83.16 89.73 +6.57 76.76 85.44 +8.86
$\checkmark$ $\times$ 70.23 76.91 +6.68 81.00 86.82 +5.82 68.81 76.32 +7.51
$\checkmark$ $\checkmark$ 69.09 75.23 +6.14 79.32 84.42 +5.10 61.09 67.95 +6.86
@@ 6 Conclusion
In this paper, we define a general paradigm for the error-based MIAs for diffusion models. Under this general paradigm, we find that the current attacks ignore the intrinsic deficiency of the diffusion model in handling the high-frequency components, which results in limited attack performance. To address this, we introduce a simple and efficient method which mitigates the negative impact of high-frequency deficiency on MIAs by filtering images’ high-frequency information. Experimental results reveal that our method can be seamlessly incorporated into attacks within the general paradigm, significantly enhancing attack performance across diverse settings.
@@
Appendix A More Details for Related Work
A.1 Denoising Diffusion Probabilistic Model (DDPM)
DDPM (Ho et al., 2020) consists of a forward process and a denoising process. In the forward process, DDPM transitions from an intractable data distribution, represented as $x_{0}\sim q_{0}(x_{0})$, to a Gaussian distribution $q_{T}(x_{T})\sim\mathcal{N}(x_{T};0,I)$. This transition is achieved by progressively adding Gaussian noise to the original image $x_{0}$. Consequently, the transition distribution at timestep $t$ is defined as follows:
@@ A.1 Denoising Diffusion Probabilistic Model (DDPM)
where $\epsilon_{\theta}(x_{t},t)$ is predicted by the diffusion models. The denoising (or reverse) process shares the same functional form as the forward process (Sohl-Dickstein et al., 2015). It is expressed as a Gaussian transition characterized by a learned mean and a fixed variance:
@@
A.2 Defense Against Model Memorization.
Wen et al. (2024) identified that when a model exactly memorizes training data, the noise prediction network exhibits a pronounced discrepancy between its conditional and unconditional predictions. Given a training data $x$, and the caption embedding $e$ consisting of $N$ tokens, they formulate the minimization objective at step $t$ as:
$\mathcal{L}(x_{t},e)=\left\|\epsilon_{\theta}(x_{t},e)-\epsilon_{\theta}(x_{t},\varnothing)\right\|_{2},$ (A.6)
where $\epsilon_{\theta}$ denotes the noise predictor and $\varnothing$ is the null prompt. The memorization score for each token at position $i\in[0,N-1]$ is then defined as:
$SS_{e_{i}}=\frac{1}{T}\sum_{t=1}^{T}\left\|\nabla_{e_{i}}\mathcal{L}(x_{t},e)\right\|_{2}.$ (A.7)
To mitigate memorization, they propose excluding a sample from the mini-batch whenever this score exceeds a predefined threshold, thereby skipping the loss computation for that sample. Since such samples have already been seen by the model during training, their removal is unlikely to degrade overall model performance. This method is proven to significantly alleviate exact memorization of training samples, providing protection for the privacy of the training set.
@@ Appendix B General Paradigm
Naive (Matsumoto et al., 2023) determines member data through the loss function, specifically judging based on the distance between the added noise and the predicted noise. SecMI (Duan et al., 2023) leverages diffusion inversion to obtain the ground truth at step $t$ and step t1t-1. Based on this, the model predicts $x_{t-1}$ from the ground truth at step $t$ and assesses whether a sample is a member by examining the distance between the predicted $x_{t-1}$ and its ground truth. PIA (Kong et al., 2023) retrieves an initial noise through a proximal initialization process, then uses the model to predict noise for samples containing that initial noise. Finally, it evaluates the membership status based on the distance between the predicted noise and the initial noise. In this section, we will demonstrate that the baselines Naive, PIA and SecMI can be translated into the general paradigm we have defined, and here we use the $\ell_{1}$ norm as an example.
@@
$\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\,t\right)||<c,$ (B.2)
@@
$\sqrt{1-\bar{\alpha}_{t}}\epsilon-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\,t\right)||<\sqrt{1-\bar{\alpha}_{t}}c,$ (B.3)
@@
$\sqrt{1-\bar{\alpha}_{t}}\epsilon-x_{t}+x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\,t\right)||<\sqrt{1-\bar{\alpha}_{t}}c.$ (B.4)
@@
$\frac{\sqrt{1-\bar{\alpha}_{t}}\epsilon-x_{t}+x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon\,,\,t\right)}{\sqrt{\bar{\alpha_{t}}}}||<\frac{\sqrt{1-\bar{\alpha}_{t}}c}{\sqrt{\bar{\alpha_{t}}}}.$ (B.5)
@@
$\frac{\sqrt{1-\bar{\alpha}_{t}}\epsilon(x_{0},0)-x_{t}+x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon(x_{0},0)\,,\,t\right)}{\sqrt{\bar{\alpha_{t}}}}||<\frac{\sqrt{1-\bar{\alpha}_{t}}c}{\sqrt{\bar{\alpha_{t}}}}.$ (B.9)
@@
SecMI: Inspired by recent works on deterministic reversing and sampling from diffusion models (Song et al., 2020b; Kim et al., 2022), SecMI used DDIM and DDIM inversion deterministic sampling in the forward and backward processes for the samples to be tested:
@@
Denote by $\bm{\Phi}_{\theta}(x_{s},t)$ the deterministic reverse, i.e., from $x_{s}$ to $x_{t}$ ($s<t$):
$x_{t}=\bm{\Phi}_{\theta}(x_{s},t)=\phi_{\theta}(\cdots\phi_{\theta}(\phi_{\theta}(x_{s},s),s+1),t-1),$ (B.14)
and $\bm{\Psi}_{\theta}(x_{t},s)$ the deterministic denoise process, i.e., from $x_{t}$ to $x_{s}$:
$x_{s}=\bm{\Psi}_{\theta}(x_{t},s)=\psi_{\theta}(\cdots\psi_{\theta}(\psi_{\theta}(x_{t},t),t-1),s+1).$ (B.15)
@@
where $\tilde{x}_{t}=\bm{\Phi}_{\theta}(x_{0},t)$. A series of transformations are designed to fetch $x_{t+1}^{target}$ and $x_{t}^{target}$ and predict $x_{t}$ by $x_{t+1}^{target}$. Therefore, Eq. B.16 can be expressed as:
$x_{t}-x_{t}^{target}||\leq\tau,$ (B.17)
@@
Appendix C Proof of Proposition 1
Proposition 1.
Denote the original standard deviations of membership scores in member and hold-out data as $\sigma_{M}$ and $\sigma_{H}$, and the standard deviations after removing the high-frequency components are $\sigma_{M}^{\prime}$ and $\sigma_{H}^{\prime}$. The standard deviation of membership scores in the high-frequency components is $h_{M}(h_{H})$, and the low-frequency components is $l_{M}(l_{H})$ in member(hold-out) data. Let $l_{H}-l_{M}=\Delta$, $h_{M}=k\cdot h_{H}$ with $k>0$. If $k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}})$, we have:
$\sigma_{H}^{\prime}/\sigma_{M}^{\prime}>\sigma_{H}/\sigma_{M}.$ (C.1)